OpenStudy (anonymous):

The 1st,3rd and 7th terms of an AP with common difference 2 are the first 3 terms of a GP.Find; a) the common ratio of the GP ganeshie8

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

say, the first term of AP is \(a\)

ganeshie8 (ganeshie8):

you're given common difference, d = 2 can you find the expressions for 3rd and 7th terms ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

after that

ganeshie8 (ganeshie8):

first find them

OpenStudy (anonymous):

3a+16

ganeshie8 (ganeshie8):

whats that ?

ganeshie8 (ganeshie8):

first term = \(a\) common difference = \(2\) 3rd term = ? 7th term = ?

OpenStudy (anonymous):

3rd term = a + 2(2) 7th term = a+6(2)

ganeshie8 (ganeshie8):

perfect ! so these 3 terms are in GP : \[\large a,~ a+4, ~a+12\]

OpenStudy (anonymous):

yes after that

ganeshie8 (ganeshie8):

find \(a\) using geometric mean : \(\large (a+4)^2 = a(a+12)\)

ganeshie8 (ganeshie8):

can you solve \(a\) in above equation ?

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

why you did (a+4)2=a(a+12) this can ypu please explain

ganeshie8 (ganeshie8):

good question :) in GP, ratio of adjacent terms is equal, right ?

ganeshie8 (ganeshie8):

\[\large a,~ a+4, ~a+12\] \(\large \dfrac{a+4}{a} = \dfrac{a+12}{a+4}\) cross multiply and solve \(a\)

OpenStudy (anonymous):

i know in gp ratio of adjacent terms is equalbut why (a+4)2=a(a+12) but why (a+4)2=a(a+12)

OpenStudy (anonymous):

wait a minute i am getting it

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

but a+4a=a+12a+4 is AP why you used gp

ganeshie8 (ganeshie8):

read the question again, the 1st, 3rd, 7th terms of given AP form first 3 terms of new GP right ?

OpenStudy (anonymous):

yes right , did you solved this before

ganeshie8 (ganeshie8):

not yet, im waiting for you to finish solving it :) let me know once you have the value of a, and the common ratio

OpenStudy (anonymous):

i solved it , a=4 r= 2

ganeshie8 (ganeshie8):

looks good ^^

ganeshie8 (ganeshie8):

good job !!

OpenStudy (anonymous):

Thanks genius man

OpenStudy (anonymous):

may god make you extreme genuis